random samples with sise 4 are drawn from the population containing the values 14, 19, 26, 31, and 53 


  1. construct a sampling distribution of the sample means.
  2. find the mean of the sample means.
  3. compute the standard error of the sample means.
1
Expert's answer
2022-04-08T03:51:13-0400

We have population values "14, 19, 26, 31, 48, 53" population size "N=6" and sample size "n=4."

Thus, the number of possible samples which can be drawn without replacement is "\dbinom{6}{4}=15."



"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Sample\ values & Sample\ mean(\bar{X}) \\ \hline
 14, 19, 26, 31 & 22.5\\
 \hdashline
 14, 19, 26, 48 & 26.75\\
 \hdashline
 14, 19, 26, 53 & 28\\
 \hdashline
 14, 19, 31, 48 & 28\\
 \hdashline
 14, 19, 31, 53 & 29.25\\
 \hdashline
 14, 19, 48, 53 & 33.5\\
 \hdashline
14, 26, 31, 48 & 29.75\\
 \hdashline
14, 26, 31, 53 & 31\\
 \hdashline
14, 26, 48, 53 & 35.25\\
 \hdashline
14, 31, 48, 53 & 36.5\\
 \hdashline
19, 26, 31, 48 & 31\\
 \hdashline
19, 26, 31, 53 & 32.25\\
 \hdashline
19, 26, 48, 53 & 36.5\\
 \hdashline
19, 31, 48, 53 & 37.75\\
 \hdashline
26, 31, 48, 53 & 39.5\\
 \hdashline
\end{array}"

a.

The sampling distribution of the sample mean "\bar{X}" is



"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 & \bar{X} & f & f(\bar{X}) & Xf(\bar{X})& X^2f(\bar{X}) \\ \hline
 & 22.5 & 1 & 1/15 & 90/60 & 8100/240\\
 \hdashline
 & 26.75 & 1 & 1/15 & 107/60 & 11449/240 \\
 \hdashline
 & 28 & 2 & 2/15 & 224/60 & 25088/240\\
 \hdashline
 & 29.25 & 1 & 1/15 & 117/60 & 13689/240 \\
 \hdashline
 & 29.75 & 1 & 1/15 & 119/60& 14161/240 \\
 \hdashline
 & 31 & 2 & 2/15 & 248/60 & 30752/240 \\
 \hdashline
 & 32.25 & 1 & 1/15 & 129/60 & 16641/240\\
 \hdashline
 & 33.5 & 1 & 1/15 & 134/60 & 17956/240 \\
 \hdashline
 & 35.25 & 1 & 1/15 & 141/60 & 19881/240 \\
 \hdashline
 & 36.5 & 2 & 2/15 & 292/60 & 42632/240 \\
 \hdashline
 & 37.75 & 1 & 1/15 & 151/60 & 22801/240 \\
 \hdashline
 & 39.5 & 1 & 1/15 & 158/60 & 24964/240 \\
 \hdashline
 Sum= & & 16 & 1 & 191/6 & 124057/120\\
 \hdashline
\end{array}"

b.The mean of the sample means is



"\mu_{\bar{X}}=E(\bar{X})=191/6=\mu"

c.



"Var(\bar{X})=\sigma^2_{\bar{X}}=E(\bar{X}^2)-(E(\bar{X}))^2""=\dfrac{124057}{120}-(\dfrac{191}{6})^2=\dfrac{7361}{360}"

The standard error of the mean



"\sigma_{\bar{X}}=\sqrt{\sigma^2_{\bar{X}}}=\sqrt{\dfrac{7361}{360}}\approx4.52186"





































"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Sample\ values & Sample\ mean(\bar{X}) \\ \hline
 14, 19, 26, 31 & 22.5\\
 \hdashline
 14, 19, 26, 48 & 26.75\\
 \hdashline
 14, 19, 26, 53 & 28\\
 \hdashline
 14, 19, 31, 48 & 28\\
 \hdashline
 14, 19, 31, 53 & 29.25\\
 \hdashline
 14, 19, 48, 53 & 33.5\\
 \hdashline
14, 26, 31, 48 & 29.75\\
 \hdashline
14, 26, 31, 53 & 31\\
 \hdashline
14, 26, 48, 53 & 35.25\\
 \hdashline
14, 31, 48, 53 & 36.5\\
 \hdashline
19, 26, 31, 48 & 31\\
 \hdashline
19, 26, 31, 53 & 32.25\\
 \hdashline
19, 26, 48, 53 & 36.5\\
 \hdashline
19, 31, 48, 53 & 37.75\\
 \hdashline
26, 31, 48, 53 & 39.5\\
 \hdashline
\end{array}"